OpenStudy (anonymous):

Select the correct difference. 5d^2 + 4d - 3 less 3d^2 - 2d + 4 2d^2 + 2d + 1 2d^2 + 6d - 7 2d^2 + 6d + 7 -2d^2 - 6d + 1

OpenStudy (anonymous):

5d^2 + 4d - 3 - (3d^2 - 2d + 4) -- distribute through the parenthesis 5d^2 + 4d - 3 - 3d^2 + 2d - 4 now just combine like terms 5d^2 - 3d^2 = 2d^2 4d + 2d = 6d -3-4 = -7 putting it all together and we have : 2d^2 + 6d - 7